Mary is playing a game in which she rolls one die and spins a spinner. What is the probability she will get both the 3 and black she needs to win the game?


1
Expert's answer
2021-01-29T03:03:52-0500

Let m denote the total number of outcomes rolling a die, therefore the probability of getting "3" is:

"P(3 \ on \ die)=\frac{1}{m}"

Let n denote the total number of outcomes spinning a spinner, therefore the probability of getting "black" is:

"P(black)=\frac{1}{n}"

Thus the probability to get both the 3 and black:

"P(3\ and\ black)=\frac{1}{m}\cdot \frac{1}{n}=\frac{1}{mn}"


Answer: "\frac{1}{mn}"